OpenStudy (anonymous):

Medals will be rewarded if anyone can help. x^6/x^?=x^4. Solve.

OpenStudy (anonymous):

Since you fanned me just type anything and I'll give you a medal.

hartnn (hartnn):

here, you can use the exponent law that, \(\Large \dfrac{x^m}{x^n}= x^{m-n}\)

hartnn (hartnn):

so in your case, m = 6, m-n = 4 just find n

hartnn (hartnn):

what did you get 'n' as ? :)

OpenStudy (anonymous):

Yeah, It's x ^2. Thanks.

hartnn (hartnn):

good! :) welcome ^_^